# Algorithmic idealism: what should you believe to experience next?

### v1

@@ Abstract
I argue for an approach to the Foundations of Physics that puts the question in the title center stage, rather than asking “what is the case in the world?”. This approach, Algorithmic Idealism, attempts to give a mathematically rigorous in-principle-answer to this question both in the usual empirical regime of physics and more exotic regimes of cosmology, philosophy, and science fiction (but soon perhaps real) technology. I begin by arguing that quantum theory, in its actual practice and in some interpretations, should be understood as telling an agent what they should expect to observe next (rather than what is the case), and that the difficulty of answering this former question from the usual “external” perspective is at the heart of persistent enigmas such as the Boltzmann brain problem, extended Wigner’s friend scenarios, Parfit’s teletransportation paradox, or our understanding of the simulation hypothesis. Algorithmic Idealism is a conceptual framework, based on two postulates that admit several possible mathematical formalizations, cast in the language of algorithmic information theory. Here I give a non-technical description of this view and show how it dissolves the aforementioned enigmas: for example, it claims that you should never bet on being a Boltzmann brain regardless of how many there are, that shutting down computer simulations does not generally terminate its inhabitants, and it predicts the apparent embedding into an objective external world as an approximate description.
@@ I Prologue: when the world is not enough
You are terminally ill. You only have a few days to live, at least that’s what the Doctor says. And yet this realization, in the bright moments between the effects of the morphine and drifting off to sleep, does not fill you with despair: you have taken precautions. Two years ago you signed a contract with AfterMath Ltd.: shortly after your death you will be scanned by a team of specially trained neuroscientists with a particular machine, and your body will be destroyed in the process. A few days later, you will be digitally resurrected in a simulated world on a computer.
@@ II Quantum probabilities as objective degrees of epistemic justification
We usually think that physics is the science of what the world is like. The idea that the contents of our physical theories represent actual things in the world, and that these things have properties that exist objectively and independently of any observer, is a characteristic assumption of many versions of scientific realism [3]. And there are good reasons to be a scientific realist: for example, realism allows us to understand that the success of science is not a miracle, but due to the fact that our theories are at least approximately representative of the truth. Intuitively, rejecting realism seems dangerously close to irrationality and pseudoscience [4]. And yet, quantum physics challenges at least the most naive versions of realism.
@@
This is evident in the scientific practice of quantum physics. Imagine a particle that has been prepared in a superposition state of two paths in an interferometer. When we measure, we will find the particle in one of the branches, and not in the other. But what is the case before the measurement? A naive stipulation that the particle is in one of the branches, but we do not know which one, leads easily to contradictions with the actual statistics we observe in experiments.
@@
We can certainly write down a wavefunction that describes all there is to say about the particle before the measurement, but as soon as we claim that the wavefunction represents “what is actually the case”, there appears a strange disconnect between the facts that we actually see (the concrete “classical” measurement outcomes) and those before we look (the quantum states). Wavefunctions do not satisfy our basic intuitions about how “facts of the world” should behave. For example, it is provably impossible to observe them directly [5]. If we think of them as real entities, then they would seem to collapse nonlocally and instantaneously upon measurement [6].
@@
Methodologically, this implies that quantum physics does not allow us to rely on our intuitions shaped by naive realism about how things in the world and their properties should behave. Now, the quantum foundations community has rightly pointed out that a lack of intuition, or of imagination, is insufficient to draw any direct metaphysical conclusions. Many phenomena that seem genuinely quantum, such as interference, entanglement, no-cloning, or teleportation, can be reproduced “classically”, i.e. within classical probability theory [7, 8, 9]. What is needed are no-go theorems that tell us that it is provably impossible to reproduce a phenomenon classically — but those results exist, and Bell’s theorem [10] is the most paradigmatic example of this. It implies that the correlations between spacelike separated events cannot have been mediated by a common cause which simply carries the predetermined value of a classical variable to the two locations, if the statistics violates a Bell inequality. Hence, our realist intuition that the local measurements simply uncover objective, preexisting facts of the world must go wrong, unless locality is violated. In scientific practice, it is then often a methodological move to give up on realist intuitions for quantum systems “in between” measurements.
@@
The difficulty of saying what quantum states tell us about the world invites us hence to shift the focus from the standard realist view towards a more empiricist perspective. We can focus on what quantum states definitely do tell us, according to the consensus of the practicing physicists: they tell us what we should expect to observe in an experiment. Given a specification of an experiment which implies a description of the involved quantum states and measurement operators, the Born rule tells us which probabilities to assign to the possible measurement outcomes. In a double-slit experiment, for example, it would be wrong to think that the quantum state tells us through which slit the particle is actually passing; but it tells us what to expect about the place at which we will observe the particle hitting the screen. In other words, Quantum Theory is about what we should believe to experience next.
@@
This methodological diagnosis is compatible with many interpretations of Quantum Theory [^1] This diagnosis is also compatible with QBism, even if the QBists would not like to phrase it in these terms. They would emphasize that the Born rule is a mere consistency requirement [30] between an agent’s personal assignments of the quantum state, the measurement operator, and the outcome probability. But this implies that any agent that has made up its mind about what quantum state and measurement operator to assign should hold beliefs about outcome probabilities that are given by the Born rule., since they all agree on how to practically use Quantum Theory successfully. But we can go one step further, and claim that this is actually what the quantum state is. This is the content of Berghofer’s degrees of epistemic justification interpretation (DEJI) [2]: “In short, DEJI is an agent-centered interpretation that views quantum mechanics as a single-user theory that allows an experiencing subject to answer the following question: Based on my experiential input, what should I believe to experience next? The input is the wave function and the output is quantum probabilities.”
@@
Berghofer’s interpretation can be understood as a modification of QBism [18] (formerly known as Quantum Bayesianism), which views quantum states as subjective degrees of belief. However, given that quantum mechanics is our most successful and fundamental theory of physics, this begs the question of “how objectivity could enter science” in QBism [2]. According to DEJI, quantum probabilities are objective degrees of epistemic justification. In this sense, quantum probabilities tell us something objective about the world: not what is the case in the world, but what we should expect Nature to answer if we ask it a question.
@@ III Restriction A: physics does not always tell agents what they should believe
It is not necessary to agree with any one of these interpretations to share the broadly empiricist sentiment that the question of “What will I see next?” is an epistemically more modest and perhaps more fruitful one than “What is really the case in the world?”. After all, the world might be a much weirder place than we have ever imagined, and humans have a bad track record of guessing the counterintuitive metaphysics of modern physics a priori. Methodologically, it makes sense to concentrate on the requirement that physics should allow us, at the very least, to make successful predictions. This includes predictions that can be intersubjectively verified, but also private ones that are in some sense irreducibly relative to an observer.
@@
A typical physicalist reaction to this question would be to deny that this is a meaningful distinction to be made in the first place: after all, there is no “fact of the world” that could ground the answer this question. The disadvantage of this reaction is that it robs us of any hope for guidance if we actually face this situation (like you did in Section I), which might well be the case within the next century or so. Moreover, according to the (either pragmatic or metaphysical) interpretation of the quantum state discussed above, the answers to questions such as “Will I really hear a detector click?” in rather mundane laboratory situations are not always grounded in facts of the world in the usual sense either [^4] I am not denying that it can be possible to learn, for example, notes in a laboratory notebook (which are also facts of the world) that tell us which quantum state the experimenter has prepared, and from this use the Born rule to obtain probabilities for our future observations. However, this procedure would be different from what I would like to call the “standard procedure”: to begin by obtaining beliefs about future facts of the world, and to deduce from this probabilities for one’s own future observations by marginalization. The reason why Restriction A is phrased in this technical form (and even more so in the publication where we introduce it in detail [17]) is to obtain a rigorous formalization of this idea of a “standard procedure”. In particular, without this qualification of what is meant in detail here, my comment would even deny the possibility of something like Algorithmic Idealism..
Indeed, recent no-go theorems tell us that we cannot always ground our beliefs about our own future observations on expected objective properties of the world, unless we give up important principles such as locality. Consider the thought experiment sketched in Figure 1 by Bong et al. [47], which is a Wigner’s friend-type [48] scenario, but extended to involve four observers and an entangled quantum state. An external observer can repeat the experiment many times, and record the outcomes (“experiences”) $a$ and $b$ of the agents Alice and Bob, given their choices of settings $x$ and $y$. Doing so, this observer may determine a relative frequency estimate of the probabilities $P(a,b|x,y)$, and compare the result with the quantum predictions. Assuming empirical adequacy of Quantum Theory, Alice and Bob can use these quantum predictions to determine what they should believe to experience in this experiment: Alice should assign the probability
@@
In Quantum Theory, we cannot always assign joint probability distributions to all variables that may, counterfactually, potentially be observed in the future, as Bell’s theorem demonstrates. However, we could have hoped that there is still a “classical” regime of the world where we can talk about variables in the usual sense that are jointly distributed, if we restrict our consideration to variables that are actually observed. Indeed, this is typically the case in our every-day world: all humans share a common Heisenberg cut, and everything that we observe during our lifetimes can (so far) be shared with all other humans. All observations (outcomes, experiences) are part of a large Boolean algebra of propositions, and we can think of a joint probability distribution that corresponds to what any hypothetical external observer should believe about the world, including the variables $a$ and $b$ from above, but astronomically many others too. What we ought to believe about our future observations should then be derived from this as a marginal probability distribution, at least in principle (in practice we certainly use heuristic simplifications and shortcuts).
However, the thought experiment above tells us that this is impossible if we also consider Charlie and Debbie. In particular, if the observed statistics $P(a,b|x,y)$ violates a so-called local friendliness inequality, then there cannot for all $x,y$ exist a joint probability distribution $P(a,b,c,d|x,y)$ of the experiences of all four observers from which $P(a,b|x,y)$ derives, unless either the principle of locality or of no-superdeterminism (or both) are violated. But if there is no such distribution, then the four observers cannot all use the “standard methodology” to determine what they should expect to experience in the experiment. That is, the four observers cannot start with what they believe about the facts of the world (including $a,b,c,d$), and from this deduce their own small part of this (the marginal distribution on $a$ and $b$). Depending on one’s interpretation of Quantum Theory, one might conclude that Charlie and Debbie can individually use the Born rule to obtain probabilities $P(c)$ and $P(d)$, but the validity of these assignments can never be verified externally.
@@
Restriction A (informally): Our physical theories do not (and sometimes cannot) give us joint predictions for the future observations of all agents.
More formally, Restriction A is a property of any given theory $T$ (such as Quantum Theory, supplemented by assumptions of locality [^5] The locality assumption of Bong et al.’s scenario [47] is weaker than the assumption of a local hidden-variable model: it only assumes that “the probability of an observable event is unchanged by conditioning on a space-like-separated free choice $z$, even if it is already conditioned on other events not in the future light-cone of $z$.” and no-superdeterminism), and predictions are assumed to be formalized as probability distributions. Restriction A may apply to some scenarios (models, experiments) within $T$, but not others. It applies to $N=4$ observers in the Wigner’s friend scenario above, and it applies to a single observer, i.e. $N=1$, in the simulation scenario of Section I.
The cases of $N=1$ and $N\geq 2$ have quite different interpretations. For $N=1$ agent, Restriction A points at a deficiency of the given physical theory: the agent might want to predict its own future observations, but cannot. For $N\geq 2$ agents, Restriction A says that the agent cannot obtain such predictions by predicting properties of the external world that are facts for all agents.
Moreover, suppose that the theory $T$ is intersubjectively empirically complete in the following sense: for all experiments for which the results can be recorded and shared among all external observers (not necessarily among all participants in the experiments, such as Wigner’s friends in boxes), the theory $T$ actually supplies us with a statistical prediction for this experiment. If this is the case, then Restriction A only applies to theory $T$ in scenarios where the agent’s observations cannot be externally recorded; in other words, where there is some sort of “epistemic horizon” [24, 25, 26] that separates some of the observers from some others. This is indeed the case in Bong et al.’s thought experiment: as we have already seen, it is impossible to record the outcomes $a,b,c,d$ in every single run of the experiment, and to obtain statistics that can, say, be published in a scientific journal. The problem is that asking Charlie or Debbie for their outcomes will in general disturb the experiment, and will prevent the violation of a local friendliness inequality. Similarly, in the simulation scenario of Section I, it was impossible for any third person (such as the Doctor) to even decide whether the event that the agent was wondering about has actually happened or not in any single implementation. After all, this was the reason for the Doctor to dismiss the patient’s question: a physicalist might want to claim that the third-person facts of the world are all there is to say, and the main characteristic of a “fact” is its intersubjective external accessibility.

### v2

@@ Abstract
I argue for an approach to the Foundations of Physics that puts the question in the title center stage, rather than asking “what is the case in the world?”. This approach, algorithmic idealism, attempts to give a mathematically rigorous in-principle-answer to this question both in the usual empirical regime of physics and in some more exotic regimes within cosmology, philosophy, and science fiction (but soon perhaps real) technology. I begin by arguing that quantum theory, in its actual practice and in some interpretations, should be understood as telling an agent what they should expect to observe next (rather than what is the case), and that the difficulty of answering this former question from the usual “external” perspective is at the heart of persistent enigmas such as the Boltzmann brain problem, extended Wigner’s friend scenarios, Parfit’s teletransportation paradox, or our understanding of the simulation hypothesis. Algorithmic idealism is a conceptual framework, based on two postulates that admit several possible mathematical formalizations, cast in the language of algorithmic information theory. Here I give a non-technical description of this view and show how it dissolves the aforementioned enigmas: for example, it claims that you should never bet on being a Boltzmann brain regardless of how many there are, that shutting down computer simulations does not generally terminate its inhabitants, and it predicts the apparent embedding into an objective external world as an approximate description.
@@ I Prologue: when the world is not enough
You are terminally ill. You only have a few days to live, at least that’s what the Doctor says. And yet this realization, in the bright moments between the effects of the morphine and drifting off to sleep, does not fill you with despair: you have taken precautions. Two years ago you signed a contract with AfterMath Ltd.: shortly after your death you will be scanned [^1] This story and most of what follows assumes that agents (at least the ones we mean here, which is ultimately supposed to include humans) can in principle be described classically, and potentially with a finite amount of classical information. This is compatible with the fact that some biological processes are genuinely quantum-mechanical and / or continuous, but a possible intuition is that these details must ultimately be irrelevant for all that matters for us as persons (including conscious experience), since robust functioning seems to forbid dependence on too fine-grained details. Whether this assumption is well-justified is the subject of contemporary debate, with a vast body of literature on e.g. the (un)suitability of “mind uploading”. For a recent proposal against this assumption, claiming that quantum states are at the heart of consciousness, see e.g. [4]. by a team of specially trained neuroscientists with a particular machine, and your body will be destroyed in the process. A few days later, you will be digitally resurrected in a simulated world on a computer.
@@
Please note that this is a summary and motivation of algorithmic idealism, intentionally dropping many details. For the precise mathematical derivations and conceptual definitions, please see the publications [2, 3].
@@ II Quantum probabilities as objective degrees of epistemic justification
We usually think that physics is the science of what the world is like. The idea that the contents of our physical theories represent actual things in the world, and that these things have properties that exist objectively and independently of any observer, is a characteristic assumption of many versions of scientific realism [6]. And there are good reasons to be a scientific realist: for example, realism allows us to understand that the success of science is not a miracle, but due to the fact that our theories are at least approximately representative of the truth. Intuitively, rejecting realism seems dangerously close to irrationality and pseudoscience [7]. And yet, quantum physics challenges [^2] To avoid potential misunderstandings from the outset, algorithmic idealism is not meant to be a new interpretation of quantum theory. It is a theory that aims to make concrete predictions in regimes where they are currently lacking (“private experiments” as discussed in Section III). In order to do so, it turns out necessary to set standard realist assumptions aside and focus exclusively on the question of the title of this paper. Quantum physics is seen as additional motivation for this focus. The goal is to construct the theory and explore its predictions (some of which are highly counterintuitive), not to have a detailed and philosophically exhaustive discussion of quantum theory, probability, or of all conceptual views that are articulated as motivation for it. at least the most naive versions of realism.
@@
This is evident in the scientific practice of quantum physics. Imagine a particle that has been prepared in a superposition state of two paths in an interferometer. When we measure, we will find the particle in one of the branches, and not in the other. But what is the case before the measurement? A naive stipulation that the particle is in one of the branches, but we do not know which one, leads easily to contradictions with the actual statistics we observe in experiments [^3] In more detail, a contradiction appears if we have a final beamsplitter that leads to perfect destructive interference, and if we naively model the experiment as a probability-$1/2$ mixture of the two experimental realizations where the particle is deterministically prepared in the upper or in the lower branch, respectively. It is simply a methodological observation that the most straightforward epistemic interpretations tend to fail to explain the experimental results. Making this simple methodological observation formally rigorous is important, but requires substantially more work, e.g. of proving that no noncontextual ontological model reproduces the specific functional form of the quantum uncertainty relations characteristic of the interference experiment [8]..
@@
We can certainly write down a wavefunction that describes all there is to say about the particle before the measurement, but as soon as we claim that the wavefunction represents “what is actually the case”, there appears a strange disconnect between the facts that we actually see (the concrete “classical” measurement outcomes) and those before we look (the quantum states). Wavefunctions do not satisfy our basic intuitions about how “facts of the world” should behave. For example, it is provably impossible to observe them directly [9]. If we think of the wave functions assigned by some observer as a real property of the physical system that it is supposed to describe, then it would seem to collapse nonlocally and instantaneously upon measurement, in potential tension with relativity (see e.g. [10, Section 6.1] for more details) [^4] These observations certainly do not rule out the possibility of realist interpretations of the quantum state. For example, Everettian interpretations evade the need for collapse, at the expense of distinguishing the absolute quantum state of the universe from the branch-relative quantum state assigned by an observer [11]. This is arguably a significant departure from our intuitive view on “facts of the world”. De Broglie-Bohm theory and objective collapse models are further examples of consistent realist interpretations, but whether they can be (or have been already) formulated in a way that avoids conflict with special relativity is subject to ongoing debate [12, 13]..
@@
Methodologically, this implies that quantum physics does not allow us to rely on our intuitions shaped by naive realism about how things in the world and their properties should behave. Now, the quantum foundations community has rightly pointed out that a lack of intuition, or of imagination, is insufficient to draw any direct metaphysical conclusions. Many phenomena that seem genuinely quantum, such as interference, entanglement, no-cloning, or teleportation, can be reproduced “classically”, i.e. by models based on classical variables that satisfy natural properties such as locality or noncontextuality [14, 15, 16]. What is needed are no-go theorems that tell us that it is provably impossible to reproduce a phenomenon classically under such assumptions — but those results exist, and Bell’s theorem [17] is the most paradigmatic example of this. It implies that the correlations between spacelike separated events cannot always be explained by a classical common cause, i.e. by a classical hidden-variable model that respects the causal structure of the experimental setup [18]. Hence, our realist intuition that the local measurements simply uncover objective, preexisting facts of the world must go wrong, unless locality is violated. In scientific practice, it is then often a methodological move to give up on realist intuitions for quantum systems “in between” measurements.
@@
The difficulty of saying what quantum states tell us about the world invites us hence to shift the focus from the standard realist view towards a more empiricist perspective. We can focus on what quantum states definitely do tell us, according to the consensus of the practicing physicists: they tell us what we should expect to observe in an experiment. Given a specification of an experiment which implies a description of the involved quantum states and measurement operators, the Born rule tells us which probabilities to assign to the possible measurement outcomes. In a double-slit experiment, for example, it would be wrong to think that the quantum state tells us through which slit the particle is actually passing; but it tells us what to expect about the place at which we will observe the particle hitting the screen. In other words, Quantum Theory is about what we should believe to observe next.
@@
This methodological diagnosis is compatible with many interpretations of Quantum Theory [^5] This diagnosis is also compatible with QBism, even if the QBists would perhaps not like to phrase it in these terms. They would emphasize that the Born rule is a mere consistency requirement [39] between an agent’s personal assignments of the quantum state, the measurement operator, and the outcome probability. But this implies that any agent that has made up its mind about what quantum state and measurement operator to assign should hold beliefs about outcome probabilities that are given by the Born rule., since they all agree on how to practically use Quantum Theory successfully. But we can go one step further, and claim that this is actually what the quantum state is. This is the content of Berghofer’s degrees of epistemic justification interpretation (DEJI) [5]: “In short, DEJI is an agent-centered interpretation that views quantum mechanics as a single-user theory that allows an experiencing subject to answer the following question: Based on my experiential input, what should I believe to experience next? The input is the wave function and the output is quantum probabilities.”
@@
Berghofer’s interpretation [^6] Berghofer’s interpretation and algorithmic idealism differ significantly: DEJI is an interpretation of quantum theory, and algorithmic idealism (AI) is not. Quantum theory matters for AI in two ways: first, as a motivation to focus on the question of what is observed next; and second, as a “litmus test” to see whether some quantum phenomena are predicted by AI. However, AI is not an interpretation of an existing theory, but a new theory that intends to make novel predictions, as outlined in the next sections. AI benefits from the conceptually rigorous formulation of DEJI, which clarifies its own use of probability substantially. can be understood as a modification of QBism [25] (formerly known as Quantum Bayesianism), which views quantum states as subjective degrees of belief. However, given that quantum mechanics is our most successful and fundamental theory of physics, this raises the question of “how objectivity could enter science” in QBism [5]. According to DEJI, quantum probabilities are objective degrees of epistemic justification. In this sense, quantum probabilities tell us something objective about the world: not what is the case in the world, but what we should expect Nature to answer if we ask it a question. In the rest of this paper, we will consider several exotic situations in which agents ask Nature a question. We are not interested in what agents believe, but what type of laws determines what they will actually probably receive as an answer — or, in other words, what they should believe to observe. Hence, Berghofer’s work will be our guiding interpretation when formulating such probabilistic laws.
@@ III Restriction A: physics does not always tell agents what they should believe
It is not necessary to agree with any one of these interpretations to share the broadly empiricist sentiment that the question of “What will I see next?” is an epistemically more modest and perhaps more fruitful one than “What is really the case in the world?”. After all, the world might be a much weirder place than we have ever imagined, and humans have a bad track record of guessing the counterintuitive metaphysics of modern physics a priori. Methodologically, it makes sense to concentrate on the requirement that physics should allow us, at the very least, to make successful predictions. This includes predictions that can be intersubjectively verified, but also private ones that are in some sense irreducibly relative to an observer [^7] Physical theories do not directly predict all intersubjectively communicable outcomes of experiments, but only those that are in some precise sense objectively definable; and similarly, algorithmic idealism does not aim for predictions of all outcomes of private experiments. For example, physics allows us to predict probabilistically whether we will hear a detector click, and algorithmic idealism says that we should be able to do the same even if the outcome is private (as in a duplication or Wigner’s friend type experiment). However, neither physics nor algorithmic idealism address directly questions of qualia, say, whether e.g. an agent will observe beauty or joy or be conscious next. We may certainly believe that these supervene on the precisely definable properties, but to understand how exactly this is the case might be so difficult to characterize that it may be a better strategy to consider the study of these question a different field of inquiry..
@@
A typical physicalist reaction to this question would be to deny that this is a meaningful distinction to be made in the first place: after all, there is no “fact of the world” that could ground the answer this question [^10] The problem is not whether it would actually be possible to scan the human being and simulate it in a functionally equivalent way — as mentioned in the first footnote of this paper, here we always make the working assumption that it is. The point is rather that there is no meaningful question to be asked from a physicalist third-person perspective, as articulated by the Doctor in Section I: the question “what will probably happen to me?” is not a question about facts of the world, as its answer is independent of any claims about e.g. how matter behaves.. The disadvantage of this reaction is that it robs us of any hope for guidance if we actually face this situation (like you did in Section I), which might well be the case within the next century or so. Furthermore, as the Boltzmann brain (BB) example below will demonstrate, there are situations in physics where we have encountered questions of a similar type (such as “should I believe that I will make a BB-type observation next?”) whose answers cannot be grounded on facts of the world either (even though it may at first seem that they could).
Indeed, recent no-go theorems tell us that we cannot always ground our beliefs about our own future observations on expected objective properties of the world, unless we give up important principles such as locality [^11] The locality assumption of Bong et al. [58] is much weaker than the locality assumption in Bell’s theorem: it does not refer to hidden variables, but expresses no-signalling for actually observed outcomes (“the probability of an observable event is unchanged by conditioning on a space-like-separated free choice $z$, even if it is already conditioned on other events not in the future light-cone of $z$.”). This is necessary to avoid conflict with relativity.. Consider the thought experiment sketched in Figure 1 by Bong et al. [58], which is a Wigner’s friend-type [59] scenario, but extended to involve four observers and an entangled quantum state. An external observer can repeat the experiment many times, and record the outcomes $a$ and $b$ of the agents Alice and Bob, given their choices of settings $x$ and $y$. Doing so, this observer may determine a relative frequency estimate of the probabilities $P(a,b|x,y)$, and compare the result with the quantum predictions. Assuming empirical adequacy of Quantum Theory, i.e. the correct prediction of the outcome probabilities of actually observed events in this scenario, Alice and Bob can use these quantum predictions to determine what they should believe to observe in this experiment: Alice should assign the probability
@@
In Quantum Theory, we cannot always assign joint probability distributions to all variables that may, counterfactually, potentially be observed in the future, as Bell’s theorem demonstrates. However, we could have hoped that there is still a “classical” regime of the world where we can talk about variables in the usual sense that are jointly distributed, if we restrict our consideration to variables that are actually observed. Indeed, this is typically the case in our every-day world: all humans share a common Heisenberg cut, and everything that we observe during our lifetimes can (so far) be shared with all other humans. All observations (outcomes, experiences) are part of a large Boolean algebra of propositions, and we can think of a joint probability distribution that corresponds to what any hypothetical external observer should believe about the world. What we ought to believe about our future observations should then be derived from this as a marginal probability distribution, at least in principle (in practice we certainly use heuristic simplifications and shortcuts).
However, this situation will change as soon as we implement extended Wigner’s friend experiments such as the one described above in our world. In this case, the results above tell us that the observations $a,b,c,d$ made by Alice, Bob, Charlie and Debbie, respectively, do not belong to a Boolean algebra of propositions of the type just described. That is, if the externally observed statistics $P(a,b|x,y)$ violates a so-called local friendliness inequality, then there cannot for all $x,y$ exist a joint probability distribution $P(a,b,c,d|x,y)$ of the observations of all four observers from which $P(a,b|x,y)$ derives, unless either the principle of locality or of no-superdeterminism (or both) are violated. But if there is no such distribution, then the four observers cannot all use the “standard methodology” to determine what they should expect to see in the experiment. That is, the four observers cannot start with what they believe about the facts of the world (including $a,b,c,d$), and from this deduce their own small part of this (the marginal distribution on $a$ and $b$). Depending on one’s interpretation of Quantum Theory, one might conclude that Charlie and Debbie can individually use the Born rule to obtain probabilities $P(c)$ and $P(d)$, but the validity of these assignments can never be verified externally.
@@
Restriction A (informally): Our physical theories do not (and sometimes cannot) give us joint predictions for the future observations obtained by all agents.
More formally, Restriction A is a property of any given theory $T$ (such as Quantum Theory, supplemented by additional assumptions such as locality and no-superdeterminism), and predictions are assumed to be formalized as probability distributions. Restriction A may apply to some scenarios (models, experiments) within $T$, but not others. It applies to $N=4$ observers in the Wigner’s friend scenario above, and it applies to a single observer, i.e. $N=1$, in the simulation scenario of Section I.
The cases of $N=1$ and $N\geq 2$ have quite different interpretations. For $N=1$ agent, Restriction A points at a deficiency of the given physical theory: the agent might want to predict its own future observations, but cannot. For $N\geq 2$ agents, Restriction A says that the agent cannot obtain such predictions by predicting properties of the external world that are facts for all agents. We will return later to the question of how to interpret this.
Moreover, suppose that the theory $T$ is intersubjectively empirically complete in the following sense: for all experiments for which the results can be recorded and shared among all external observers (not necessarily among all participants in the experiments, such as Charlie and Debbie), the theory $T$ actually supplies us with a statistical prediction for this experiment. If this is the case, then Restriction A only applies to theory $T$ in scenarios where the agent’s observations cannot be externally recorded; in other words, where there is some sort of “epistemic horizon” [33, 34, 35] that separates some of the observers from some others. This is indeed the case in Bong et al.’s thought experiment: as we have already seen, it is impossible to record the outcomes $a,b,c,d$ in every single run of the experiment, and to obtain statistics that can, say, be published in a scientific journal. The problem is that asking Charlie or Debbie for their outcomes will in general disturb the experiment, and will prevent the violation of a local friendliness inequality. Similarly, in the simulation scenario of Section I, it was impossible for any third person (such as the Doctor) to even decide whether the event that the agent was wondering about has actually happened or not in any single implementation. After all, this was the reason for the Doctor to dismiss the patient’s question: a physicalist might want to claim that the third-person facts of the world are all there is to say, and the main characteristic of a “fact” is its intersubjective external accessibility.
@@
Algorithmic idealism, as described below, can be understood as a reaction to Restriction A in the following sense. In this conceptual framework, Restriction A for any given single agent ($N=1$) is seen as a serious problem that needs to be alleviated, because it reveals the inability to predict something about the results of actual experiments that somebody can do. For example, you can in principle (and, in the perhaps not so far future, in practice) agree on being copied into a computer simulation in some sense, and see what happens to you. Despite the fact that the outcome cannot be intersubjectively shared (as explained above, this is a prerequisite for Restriction A to even apply in the first place), it seems absurd to believe that there was nothing that could be said (probabilistically, say, or plausibilistically [30], or in some other sort of mathematical formulation) about which sort of outcome of this experiment you ought to expect to experience. After all, the world will kick back at you and generate some subjective outcome, and something must determine the nature of this outcome. Even claiming that nothing can be said in this case is just an utterance of a human being in need of clarification, which means in need of mathematical formalization. Indeed, algorithmic idealism suggests a particular resolution of Restriction A for $N=1$ agents via Postulate 2 below (and via its formalization in any specifically chosen mathematical model).
However, algorithmic idealism does not claim to resolve Restriction A for $N\geq 2$ agents, but rather suggests to accept it: according to its predictions, different agents do not in general share a common world into which they seem to be embedded. Instead, “objective reality” is an approximate and emergent statistical phenomenon that does not apply in all cases (see Subsection VI.3 below). Indeed, Restriction A for $N\geq 2$ observers (and the recent results [58] showing that, under some natural assumptions, all future physical theories will suffer from it) is taken as corroborating evidence for the correctness of algorithmic idealism’s strategy to drop the assumption that agents are fundamentally embedded into external worlds.